# Concurrency Constrained Scheduling with Tree-Like Constraints

### v1

@@ 1 Introduction
Following the three-field notation of Graham et al. for scheduling problems [28], we denote the concurrency constrained setting with unit processing times by $P_{n}\mid conc,p_{j}=1\mid*$. The first field denotes the machine model being used, in our case $n$ identical machines that run in parallel. The second field denotes constraints on the problem, where ‘conc’ denotes that concurrency constrains (via a conflict graph) are present, while $p_{j}=1$ signifies that all jobs require a unit amount of processing time on any of the $n$ machines. Additional constraints and restrictions can be added to this field, as will be discussed later on. Finally, the last field ‘*’ denotes the objective function that we optimise.
@@
We denote this setting by $P_{n}\mid conc\mid*$. Note that in the case of arbitrary processing times, some proper colorings of $G$ no longer correspond to feasible schedules. However, every feasible schedule is still a proper coloring.
@@
On the negative side, we show that if either jobs have unbounded processing times or if release times are introduced, then only $P_{n}\mid conc,r_{j}\geq 0,p_{j}=1\mid C_{\max}$ remains FPT; all other problems become XALP-complete. Our hardness results here are obtained by developing reductions from the XALP-complete Precoloring Extension problem. We note that XALP-completeness results have two important implications. First, these problems are in XP, i.e., have an algorithm with running time of the form $O(|I|^{f(k)})$, where $|I|$ is the size of the instance. Second, a conjecture by Pilipczuk and Wrochna [50] implies that XALP-hard problems are unlikely to have an XP-algorithm with space usage $f(k)\cdot|I|^{O(1)}$.
@@ 3 Bounded Processing Times
In the following section we focus on cases where job processing times are either unit, or relatively small, i.e. $P_{n}\mid conc,p_{j}=1\mid*$ and $P_{n}\mid conc,p_{j}\leq p\mid*$ for some fixed $p\in O(1)$. Previous work has shown that for the $\sum C_{j}$ objective, there always exists an optimal schedule $C$ with makespan at most $tw(G)\cdot\log n$ for the unit processing time case [36], and $tw(G)\cdot p\log n$ for the $p_{j}\leq p$ case [30]. In what follows we generalize and unify these results into our framework. For this, we first introduce the notion of minimal feasible schedules.
@@
We begin by considering the case of unit processing times. Note that this case is equivalent to the case of $p_{j}=p$ for all jobs $j$, since one can simply multiply the completion times in any feasible schedule for a $P_{n}\mid conc,p_{j}=1\mid*$ instance by $p$. We will heavily rely on the notion of a Grundy number of a graph: A greedy coloring of a graph $G=(V,E)$ is a proper coloring $C:\{1,\ldots,n\}\rightarrow\mathbb{N}$ such that for every vertex $j$ and any color $c\in\{1,\ldots,C(j)-1\}$ there is some vertex $i$ with $\{i,j\}\in E$ and $C(i)=c$. The value of a greedy coloring $C$ of $G=(V,E)$ is $C_{\max}=\max_{v\in V}C(v)$. The Grundy number of $\Gamma(G)$ of a graph $G$ is the maximum value of any greedy coloring of $G$. From [56], we have:
@@
A feasible minimal schedule of any $P_{n}\mid conc,p_{j}=1\mid*$ instance corresponds to a greedy coloring of the conflict graph $G$.
@@
Consider a feasible minimal schedule for the set of jobs $\{1,\ldots,n\}$ given in some $P_{n}\mid conc,p_{j}=1\mid*$ instance, and let $j\in\{1,\ldots,n\}$. We argue that for every $c\in\{1,\ldots,C(j)-1\}$ there exists some $i\in\{1,\ldots,n\}$ with $\{i,j\}\in E$ and $C(i)=c$. Suppose not. That is, suppose there is some $c\in\{1,\ldots,C(j)-1\}$ such that for all $i\in\{1,\ldots,n\}$ with $\{i,j\}\in E$ we have $C(i)\not=c$. Then, since all processing times are unit, the schedule $C^{\prime}$ with $C^{\prime}(j)=c$ and $C^{\prime}(i)=C(i)$ for all $i\in\{1,\ldots,n\}\setminus\{j\}$ is feasible as well, contradicting our assumption that $C$ is minimal. It follows that $C$ is a greedy coloring of $G$. ∎
Recall that a scheduling criterion is regular if it is non-decreasing in the completion time of the jobs; i.e., decreasing the completion time of any job does not decrease the value of the schedule. We thus combine Lemma 1 and 2 to get:
@@
Any $P_{n}\mid conc,p_{j}=1\mid*$ instance where * is regular has an optimal schedule $C$ with $C_{\max}\leq tw\cdot\log n+1$.
@@
We next consider the case where the processing times of jobs may vary, but are still rather small. We denote this case by $P_{n}\mid conc,p_{j}\leq p\mid*$. We extend the bound of Corollary 2 to this case as well, exploiting the connection between treedepth and treewidth, although at the price of being slightly less precise.
@@
There exists a feasible minimal schedule $C$ for any $P_{n}\mid conc,p_{j}\leq p\mid*$ instance such that $C_{\max}=O(tw(G)\cdot p\log n)$.
@@ 4 Unbounded Processing Times
We next consider the case where the processing times of jobs are unbounded, i.e. problems of the form $P_{n}\mid conc\mid*$. We show that already for the two most basic objective functions of $C_{\max}$ and $\sum C_{j}$, the corresponding scheduling with concurrency constrained problem becomes XALP-complete parameterized by the treewidth of the conflict graph. As $P_{n}\mid conc\mid C_{\max}$ and $P_{n}\mid conc\mid\sum C_{j}$ are special cases of any $P_{n}\mid conc\mid*$ problem considered in this paper, this will complete the picture for the second row in Table 1. For both these problems, we show XALP-hardness by a reduction from Precoloring Extension.
@@ 6 Discussion
We may also consider further aspects of scheduling problems. First, if we consider preemption, then $p_{j}\geq 1$ but the times at which a job is processed is not necessarily a contiguous interval. The problem $P_{n}\mid conc,pmtn\mid C_{\max}$ is known as Preemptive Multicoloring, Multichromatic Number, Weighted Coloring, or Minimum Integer Weighted Coloring [3, 4, 29, 32, 33, 34, 48, 57]. We can reduce Preemptive Multicoloring to Coloring by replacing each job $j$ by $p_{j}$ jobs that are pairwise in conflict. That is, we replace each vertex by $p_{j}$ (true) twins. Hence, if $\max_{j}p_{j}$ is bounded or the $p_{j}$ are given in unary, all known algorithmic results automatically apply for graph classes that are closed under adding true twins (e.g., all hereditary graph classes that do not forbid small cliques, graphs of bounded treewidth). We thus ask what happens when the $p_{j}$ are in binary and not bounded. We may also consider $P_{n}\mid conc,pmtn\mid\sum C_{j}$, known as Preemptive Multichromatic Sum [6, 30].
@@ Acknowledgments
The first author wants to thank Meike Hatzel, Pascal Gollin, Matjaž Krnc, and Martin Milanič for discussions on Chromatic Sum and related problems.

### v2

@@ 1 Introduction
Following the three-field notation of Graham et al. for scheduling problems [28], we denote the concurrency constrained setting with unit processing times by $P_{n}\mid conc,p_{j}=1\mid\ast$. The first field denotes the machine model being used, in our case $n$ identical machines that run in parallel. The second field denotes constraints on the problem, where ‘conc’ denotes that concurrency constrains (via a conflict graph) are present, while $p_{j}=1$ signifies that all jobs require a unit amount of processing time on any of the $n$ machines. Additional constraints and restrictions can be added to this field, as will be discussed later on. Finally, the last field ‘*’ denotes the objective function that we optimise.
@@
We denote this setting by $P_{n}\mid conc\mid\ast$. Note that in the case of arbitrary processing times, some proper colorings of $G$ no longer correspond to feasible schedules. However, every feasible schedule is still a proper coloring.
@@
On the negative side, we show that if jobs have unbounded processing times or release times, then only $P_{n}\mid conc,r_{j}\geq 0,p_{j}=1\mid C_{\max}$ remains FPT; all other problems become XALP-complete. Our hardness results here are obtained by developing reductions from the XALP-complete Precoloring Extension problem. We note that XALP-completeness results have two important implications. First, these problems are in XP, i.e., have an algorithm with running time of the form $O(|I|^{f(k)})$, where $|I|$ is the size of the instance. Second, a conjecture by Pilipczuk and Wrochna [50] implies that XALP-hard problems are unlikely to have an XP-algorithm with space usage $f(k)\cdot|I|^{O(1)}$.
@@ 3 Bounded Processing Times
In the following section we focus on cases where job processing times are either unit, or relatively small, i.e. $P_{n}\mid conc,p_{j}=1\mid\ast$ and $P_{n}\mid conc,p_{j}\leq p\mid\ast$ for some fixed $p\in O(1)$. Previous work has shown that for the $\sum C_{j}$ objective, there always exists an optimal schedule $C$ with makespan at most $tw(G)\cdot\log n$ for the unit processing time case [36], and $tw(G)\cdot p\log n$ for the $p_{j}\leq p$ case [30]. In what follows we generalize and unify these results into our framework. For this, we first introduce the notion of minimal feasible schedules.
@@
We begin by considering the case of unit processing times. Note that this case is equivalent to the case of $p_{j}=p$ for all jobs $j$, since one can simply multiply the completion times in any feasible schedule for a $P_{n}\mid conc,p_{j}=1\mid\ast$ instance by $p$. We will heavily rely on the notion of a Grundy number of a graph: A greedy coloring of a graph $G=(V,E)$ is a proper coloring $C:\{1,\ldots,n\}\rightarrow\mathbb{N}$ such that for every vertex $j$ and any color $c\in\{1,\ldots,C(j)-1\}$ there is some vertex $i$ with $\{i,j\}\in E$ and $C(i)=c$. The value of a greedy coloring $C$ of $G=(V,E)$ is $C_{\max}=\max_{v\in V}C(v)$. The Grundy number of $\Gamma(G)$ of a graph $G$ is the maximum value of any greedy coloring of $G$. From [56], we have:
@@
A feasible minimal schedule of any $P_{n}\mid conc,p_{j}=1\mid\ast$ instance corresponds to a greedy coloring of the conflict graph $G$.
@@
Consider a feasible minimal schedule for the set of jobs $\{1,\ldots,n\}$ given in some $P_{n}\mid conc,p_{j}=1\mid\ast$ instance, and let $j\in\{1,\ldots,n\}$. We argue that for every $c\in\{1,\ldots,C(j)-1\}$ there exists some $i\in\{1,\ldots,n\}$ with $\{i,j\}\in E$ and $C(i)=c$. Suppose not. That is, suppose there is some $c\in\{1,\ldots,C(j)-1\}$ such that for all $i\in\{1,\ldots,n\}$ with $\{i,j\}\in E$ we have $C(i)\not=c$. Then, since all processing times are unit, the schedule $C^{\prime}$ with $C^{\prime}(j)=c$ and $C^{\prime}(i)=C(i)$ for all $i\in\{1,\ldots,n\}\setminus\{j\}$ is feasible as well, contradicting our assumption that $C$ is minimal. It follows that $C$ is a greedy coloring of $G$. ∎
Recall that a scheduling criterion is regular if it is non-decreasing in the completion time of the jobs; i.e., decreasing the completion time of any job does not increase the value of the schedule. We thus combine Lemma 1 and 2 to get:
@@
Any $P_{n}\mid conc,p_{j}=1\mid\ast$ instance where * is regular has an optimal schedule $C$ with $C_{\max}\leq tw\cdot\log n+1$.
@@
We next consider the case where the processing times of jobs may vary, but are still rather small. We denote this case by $P_{n}\mid conc,p_{j}\leq p\mid\ast$. We extend the bound of Corollary 2 to this case as well, exploiting the connection between treedepth and treewidth, although at the price of being slightly less precise.
@@
There exists a feasible minimal schedule $C$ for any $P_{n}\mid conc,p_{j}\leq p\mid\ast$ instance such that $C_{\max}=O(tw(G)\cdot p\log n)$.
@@ 4 Unbounded Processing Times
We next consider the case where the processing times of jobs are unbounded, i.e. problems of the form $P_{n}\mid conc\mid\ast$. We show that already for the two most basic objective functions of $C_{\max}$ and $\sum C_{j}$, the corresponding scheduling with concurrency constrained problem becomes XALP-complete parameterized by the treewidth of the conflict graph. As $P_{n}\mid conc\mid C_{\max}$ and $P_{n}\mid conc\mid\sum C_{j}$ are special cases of any $P_{n}\mid conc\mid\ast$ problem considered in this paper, this will complete the picture for the second row in Table 1. For both these problems, we show XALP-hardness by a reduction from Precoloring Extension.
@@ 6 Discussion
We may also consider further aspects of scheduling problems. First, if we consider preemption, then $p_{j}\geq 1$ but the times at which a job is processed is not necessarily a contiguous interval. The problem $P_{n}\mid conc,pmtn\mid C_{\max}$ is known as Preemptive Multicoloring, Multichromatic Number, Weighted Coloring, or Minimum Integer Weighted Coloring [3, 4, 29, 32, 33, 34, 48, 57]. We can reduce Preemptive Multicoloring to Coloring by replacing each job $j$ by $p_{j}$ jobs that are pairwise in conflict. That is, we replace each vertex by $p_{j}$ (true) twins. Note that this step can increase the pathwidth and treewidth of the conflict graph by a multiplicative factor of $\max p_{j}$. Thus, we can transfer positive results for the non-preemptive case with $p_{j}=1$ to positive results for the preemptive case with $p_{j}\geq 1$, when taking $\max p_{j}$ as additional parameter. It is interesting to investigate what happens when the $p_{j}$ are not parameter bounded. We may also consider $P_{n}\mid conc,pmtn\mid\sum C_{j}$, known as Preemptive Multichromatic Sum [6, 30].
@@ Acknowledgments
The first author wants to thank Meike Hatzel, Pascal Gollin, Matjaž Krnc, and Martin Milanič for discussions on Chromatic Sum and related problems. An extended abstract of this paper appears in the proceedings of the 51st International Workshop on Graph-Theoretic Concepts in Computer Science, WG 2025, published in the Springer Lecture Notes in Computer Science series.